What is the magnitude of the force acting on a straight wire of length 1.029 m, with a current of 8.71 A, and inclined at an angle of 23.44 degrees to a uniform magnetic field of magnitude 0.692 T?




1
Expert's answer
2021-06-03T09:20:10-0400
"F=BIL\sin{23.44}\\F=(0.692)(8.71)(1.029)\sin{23.44}=2.47\ N"